OpenStudy (anonymous):

what is the geometric mean of 3 and 21? A) 3 square root of 7 B) 7 square root of 3 C) 31.5 D) 63

OpenStudy (anonymous):

A

hero (hero):

\[GM = \sqrt{ab}\] In this case, a = 3, b = 21

OpenStudy (radar):

\[\sqrt{3 * 21}\]\[\sqrt{63}\]\[\sqrt{9}\sqrt{7}\] \[3\sqrt{7}\]

OpenStudy (anonymous):

3 square root of seven

OpenStudy (radar):

yes